# Group-robust Machine Unlearning

### v1

@@ Abstract
Machine unlearning is an emerging paradigm to remove the influence of specific training data (i.e., the forget set) from a model while preserving its knowledge of the rest of the data (i.e., the retain set). Previous approaches assume the forget data to be uniformly distributed from all training datapoints. However, if the data to unlearn is dominant in one group, we empirically show that performance for this group degrades, leading to fairness issues. This work tackles the overlooked problem of non-uniformly distributed forget sets, which we call group-robust machine unlearning, by presenting a simple, effective strategy that mitigates the performance loss in dominant groups via sample distribution reweighting. Moreover, we present MIU (Mutual Information-aware Machine Unlearning), the first approach for group robustness in approximate machine unlearning. MIU minimizes the mutual information between model features and group information, achieving unlearning while reducing performance degradation in the dominant group of the forget set. Additionally, MIU exploits sample distribution reweighting and mutual information calibration with the original model to preserve group robustness. We conduct experiments on three datasets and show that MIU outperforms standard methods, achieving unlearning without compromising model robustness. Source code available here. ††Corresponding Author: thomas.demin@unitn.it
@@ 1 Introduction
In several countries, recent regulations grant users greater control over their digital privacy, explicitly recognizing their right to request the removal of personal data, known as the “right to be forgotten” [32, 45]. Therefore, to comply with such regulations, machine learning systems should be able to forget specific training data upon user request—a challenge addressed by machine unlearning [28, 9].
@@
In the machine unlearning literature, the standard setup assumes that a group of individuals requests the removal of their data. A machine unlearning algorithm then fulfills this request by making the model “forget” the designated data (i.e., the forget set) while maintaining its accuracy on the test set [28]. The naive solution, called exact unlearning, consists in retraining the model without the forget data, however, this is computationally prohibitive for large models [52]. Approximate unlearning methods [23, 13, 28] overcome this limitation by unlearning the model with significantly less resources but less unlearning guarantees.
@@
Most approaches assume that the forget sets are uniformly sampled from the training data [23, 13, 7, 43, 18, 52, 20]. However, studies [3, 48] show that individuals from different social and cultural backgrounds (i.e., groups) request to be forgotten at varying rates: [3] reports that 44.4% of unlearning requests to news websites relate to professional wrongdoing or crimes, while [48] finds that wealthier, highly educated individuals are more likely to request unlearning. Existing methods overlook this imbalance, potentially degrading the accuracy of dominant groups in the forget set and leading to unfair outcomes (see Figs. 1 and 2). This can be critical in scenarios requiring high accuracy across all groups. In a recommendation system, if unlearning requests predominantly come from a specific group (e.g., young males), the system’s quality for that group may degrade, potentially making it unusable. Thus, machine unlearning algorithms must account for distribution shifts.
@@
We target this unexplored scenario, which we name group-robust machine unlearning, that aims to unlearn the user’s data while minimizing the performance deterioration of groups dominating the forget set. Differently from previous works, we tackle both exact and approximate unlearning by (i) finding a retraining strategy that preserves the original group robustness [42, 30], and (ii) proposing an approximate unlearning method that efficiently forgets data while preserving robustness for unbalanced forget sets.
@@
For (i), we show that reweighting the sampling distribution during retraining compensates for information loss with minimal robustness impact. We validate this strategy (called reweight) against group-DRO [42], a popular group-robust optimization method, showing that reweight better preserves model group robustness in exact unlearning. For (ii), we introduce an approximate unlearning method called MIU (Mutual Information-aware Machine Unlearning) leveraging mutual information [2] minimization and calibration to unlearn the forget set while preserving model robustness. By minimizing the mutual information between forget-set features and ground-truth labels, we decorrelate unlearning from spurious attributes [30], mitigating performance loss for dominant groups. To prevent affecting other groups, we calibrate the unlearned model’s mutual information to match the original one. Coupled with reweight, MIU outperforms established unlearning approaches (L1-sparse [23], SalUn [13], and SCRUB [28]) on CelebA [31], Waterbirds [42], and FairFace [24] in both unlearning efficacy and preserved group robustness.111We use robustness and group-robustness interchangeably in this paper.
@@ 2 Related work
Machine unlearning. Exact machine unlearning methods [4, 46, 1] guarantee that sensitive data is removed. However, they are impractical [9, 28] as retraining (part of) the model is prohibitive [35]. Instead, approximate unlearning concentrates on computational feasibility by relaxing the guarantees constraints [28, 9, 23, 13, 5]. Most works focus on random [23, 13, 7, 43, 18, 52, 20], and class [5, 9, 10, 19, 8, 51] unlearning, respectively forgetting i.i.d. data points, and all data points from a single class. While the former lowers the forget-set accuracy to match that of the model retrained without the unlearning data, the latter aims at scoring zero accuracy on the unlearned class. A few works [9, 14, 8] explore subclass unlearning, where a subset of a class (e.g., subclass car from superclass vehicles) is removed from model weights. While subclass and group-robust unlearning share similarities, the latter focuses on preserving group accuracies, whereas subclass unlearning alters the model to behave as if the target subclass was never in the training set. Related to our research, [12] suggests a bi-level optimization to sample adversarial forget sets that are difficult to unlearn. Also related, [6] proposes machine unlearning as an efficient debiasing technique. Instead, [48] investigates how uniform and non-uniform data sampling affect fairness in MLPs and tabular data, limiting the evaluation to exact unlearning.
Compared to previous works, this paper investigates and addresses the performance loss caused by non-uniformly distributed forget datasets, proposing an exact and approximate unlearning algorithm to mitigate this issue.
@@
Group-robust learning. Methods for group-robust optimization train deep learning models to be robust to spurious correlations. Algorithms are categorized based on their access to group information. Within those that assume access to group annotations [42, 15, 49, 21, 26], group-DRO [42] dynamically reweights the misclassification penalty for each group to optimize the worst-group accuracy. Instead, Idrissi et al. [21] propose a simple baseline that subsamples each group to match the size of the smallest one. While these works usually achieve better results, accessing the group information can be challenging (e.g., annotation cost). Within methods agnostic to group information [30, 50, 21, 44, 33], JTT [30] increases the sampling probability of wrongly classified data to improve worst-group accuracy. Correct-N-Contrast [50] uses a contrastive loss to pull correctly and misclassified samples of the same class while pushing apart wrongly classified data points of different categories. However, methods that require group information also show state-of-the-art performance on groups discovered from data [25, 11].
@@ 3 Method
This section formulates the machine unlearning task (Sec. 3.1) and the group-robust machine unlearning problem (Sec. 3.2). Then, it introduces the proposed sample distribution reweighting strategy (Sec. 3.3), showing its effectiveness in group-robust machine unlearning. Finally, Sec. 3.4 describes MIU, our approximate unlearning method tailored for group-robust unlearning.
@@ 3.1 Machine unlearning
Let be a learnable function, where is a non-linear feature extractor parameterized by , and is a linear classifier parameterized by , mapping inputs from the image to the target space . Let be a training dataset of size , where is an image, and its target label (e.g., age). The machine unlearning goal is to scrub the influence of a desired forget set from the pretrained model, trained on , without altering its performance on retain set and test data . Let the original model trained on with algorithm be denoted as (or Pretrain). A successful unlearning algorithm outputs scrubbed weights , such that is as close as possible to the exact unlearning model (or Retrain), trained solely on with algorithm [28, 13, 52].
@@ 3.2 Group-robust machine unlearning
In group-robust machine unlearning, we consider the forget data non-uniformly distributed. Therefore, let us re-define the training dataset as , where and are defined as in Sec. 3.1, and is the protected or sensitive attribute (e.g., gender, ethnicity). Now, let be the set of all groups, defined as the cartesian product between the target label set and the protected attribute set [42, 26]. We denote the -th datapoint group as . Each target-sensitive attribute pair (e.g., males between the ages of 20-29) identifies a unique group.222Target and protected attributes are color-coded.
The more samples of a group are removed, the lower the resulting group accuracy. Intuitively, removing non-uniformly distributed data changes the retain set group distribution, ultimately harming the model’s generalization performance on the dominant group of the forget set. Figure 2 shows the accuracy degradation caused by removing different percentages of attractive males from the CelebA [31] dataset, using the L1-sparse [23], SalUn [13], and SCRUB [28] approaches. Further analysis is provided in the Appendix.
@@ 3.3 Frustratingly easy group-robust unlearning
Section 3.2 introduces the group-robust machine unlearning task and shows the extent to which the accuracy of the forget set dominant group(s) drops after unlearning. Unlike prior works (see Sec. 1), we aim at unlearning non-uniform forget data while preserving the model accuracy on the dominant group of the forget set. Therefore, this section proposes an exact unlearning strategy tailored to this task. To mitigate the performance degradation of the dominant group of , we argue that reweighting the data distribution to account for the removed samples is a simple and effective baseline that retrains the model with a minimal performance drop. Intuitively, increasing the sampling likelihood for partly unlearned groups rebalances the retain set group statistics to match those of the training dataset.
Formally, let be the probability of sampling , let respectively be the group frequencies of the training and retain datasets. We reweight according to the ratio : , where is the adjusted group weight for sample . We denote the data distribution reweighting strategy (or reweight) as , i.e., denotes the reweighing strategy applied to the retain dataset.
We validate reweight by comparing the model retrained with reweight and group-DRO [42]. The popularity of group-DRO [42] in group-robust optimization suggests it should minimize drops in the forget-set dominant group accuracy. Figure 3 summarizes the outcome of our analysis. As a byproduct of strongly optimizing worst-group accuracies, we notice that group-DRO [42] can also increase the forget set accuracy. This issue makes approximate unlearning evaluation more difficult if Retrain + group-DRO is used as the gold standard. Assuming a hypothetical original forget-set accuracy of 70%, if an approximate unlearning algorithm targeting such a gold standard leads to higher accuracy (e.g., 80%), then was the knowledge unlearned if forget-set accuracy increased? Answering this question is non-trivial as the retrained model cannot be accessed for real applications. Furthermore, retraining with group-DRO [42] causes a test accuracy drop (-5.8%) in FairFace [24]. On the contrary, reweight does not suffer from these issues.
@@ 3.4 Group-robust unlearning without retraining
Section 3.3 shows that sample reweighting is a simple and effective approach to machine unlearning that preserves the original model’s robustness. However, model retraining is inefficient [9, 28] and unpractical [35]. Therefore, this section proposes MIU, our approximate machine unlearning algorithm that jointly tackles unlearning while decorrelating unlearning and group robustness.
As scrubbing unlearning data affects the forget set dominant group accuracy, we propose a unified objective for jointly unlearning while preserving original group robustness. We use mutual information between output features and group annotation, which upon minimization on forget data jointly unlearns and mitigates the performance degradation on the dominant group of the forget set [40].
Formally, let be the mutual information between random variables and , associated with model features and group , then:
where is the joint pdf of and , and , are marginal pdfs of and . Minimizing reduces the dependency between and while also reducing it between and , jointly unlearning the network and disentangling features and protected attributes. However, computing the mutual information for continuous variables is generally intractable [36, 2]. Therefore, we follow [2], and estimate the mutual information with an MLP :
where refers to a sampling from the joint distribution, refers to a sampling from the product of the marginal distributions. To sample from the product of the marginal distributions we sample twice from the joint distribution: , and keep only and as in [2]. We now use this definition to derive MIU (Mutual Information-aware Machine Unlearning), our proposed method.
@@
Intuitively, this term is high when the forget features correlate with the group information. Ideally, we want this term to be low as this will imply that we are unlearning the relation between forget-set features and labels while decorrelating the unlearning process from group information.
Calibration term. As unlearning might also affect other groups, we designed an extra term to improve group performance retention. Thus, we minimize the mutual information discrepancy between the original and unlearning model on the reweighted retain dataset :
where estimates the mutual information using features computed by the original backbone . Equation 4 encourages the unlearned model to mimic the original robustness, preserving its behavior across groups.
Retaining term. Since unlearning generally causes performance degradation [28, 13, 23], we perform gradient descent steps on the retain data to ensure that the original model group accuracy is preserved after unlearning. Additionally, we use reweight to limit group robustness degradation. Therefore, we optimize the cross-entropy loss on the reweighted retain dataset :
@@
Further details are in the Appendix.

### v2

@@ Abstract
Machine unlearning is an emerging paradigm to remove the influence of specific training data (i.e., the forget set) from a model while preserving its knowledge of the rest of the data (i.e., the remaining set). Previous approaches assume the forget data to be uniformly distributed from all training datapoints. However, if the data to unlearn is dominant in one group (e.g., ethnicity, gender), we empirically show that performance for this group can degrade, leading to fairness issues. To perform unlearning while preserving fairness, this work addresses the overlooked problem of non-uniformly distributed forget sets, which we refer to as group-robust machine unlearning. We formalize the problem and present a simple and effective exact unlearning strategy that mitigates the performance loss in dominant groups via sample distribution reweighting. Moreover, we present MIU (Mutual Information-aware Machine Unlearning), the first approach for group robustness in approximate machine unlearning. MIU minimizes the mutual information between model features and group information, achieving unlearning while reducing performance degradation in the dominant group of the forget set. Additionally, MIU exploits sample distribution reweighting and mutual information calibration with the original model to preserve group robustness. We conduct experiments on three datasets and show that MIU outperforms standard methods, achieving unlearning without compromising model robustness. Source code available at https://github.com/tdemin16/group-robust_machine_unlearning.
@@ 1 Introduction
In several countries, recent regulations grant users greater control over their digital privacy, explicitly recognizing their right to request the removal of personal data, known as the “right to be forgotten” (Mantelero, 2013; Voigt & Von dem Bussche, 2017). Therefore, to comply with such regulations, machine learning systems should be able to forget specific training data upon user request—a challenge addressed by machine unlearning (Kurmanji et al., 2024; Chundawat et al., 2023a).
@@
In the machine unlearning literature, the standard setup assumes that a group of individuals requests the removal of their data. A machine unlearning algorithm then fulfills this request by making the model “forget” the designated data (i.e., the forget set) while maintaining its utility (Kurmanji et al., 2024). The naive solution, called exact unlearning, consists of retraining the model without the forget data; however, this is computationally prohibitive for large models (Zhao et al., 2024b). Approximate unlearning methods (Jia et al., 2023; Fan et al., 2024b; Kurmanji et al., 2024) overcome this limitation by unlearning the model with significantly fewer resources but with less unlearning guarantees.
@@
Most approaches assume that the forget sets are uniformly sampled from the training data (Jia et al., 2023; Fan et al., 2024b; Cheng & Amiri, 2023; Shen et al., 2024; He et al., 2024; Zhao et al., 2024b; Huang et al., 2024). However, studies (Bertram et al., 2019; Zhang et al., 2024) show that individuals from different social and cultural backgrounds (i.e., groups) request to be forgotten at varying rates: Bertram et al. (2019) reports that 44.4% of unlearning requests to news websites relate to professional wrongdoing or crimes, while Zhang et al. (2024) finds that wealthier, highly educated individuals are more likely to request unlearning. Existing methods overlook this imbalance, potentially degrading the accuracy of dominant groups in the forget set and leading to unfair outcomes (see Figs.˜1 and 2). This can be critical in scenarios requiring high accuracy across all groups. In a recommendation system, if unlearning requests predominantly come from a specific group (e.g., young males), the system’s quality for that group may degrade, potentially making it unusable. Thus, machine unlearning algorithms must account for distribution shifts.
@@
We target this unexplored scenario, which we name group-robust machine unlearning, that aims to unlearn the user’s data while minimizing the performance deterioration of groups dominating the forget set. Differently from previous works, we tackle both exact and approximate unlearning by (i) finding a retraining strategy that preserves the original group robustness (Sagawa et al., 2020), and (ii) proposing an approximate unlearning method that efficiently forgets data while preserving robustness for unbalanced forget sets.
@@
For (i), we show that reweighting the sampling distribution during retraining compensates for information loss with minimal robustness impact. We validate this strategy (called reweight) against group-DRO (Sagawa et al., 2020), a popular group-robust optimization method, showing that reweight better preserves model group robustness in exact unlearning. For (ii), we introduce an approximate unlearning method called MIU (Mutual Information-aware Machine Unlearning), leveraging mutual information minimization (Belghazi et al., 2018) and calibration to unlearn the forget set while preserving model robustness. By minimizing the mutual information between forget-set features and ground-truth group annotations, we decorrelate unlearning from spurious attributes (Liu et al., 2021), mitigating performance loss for dominant groups. To prevent affecting other groups, we calibrate the unlearned model’s mutual information to match the original one. Coupled with reweight, MIU outperforms established unlearning approaches (L1-sparse (Jia et al., 2023), SalUn (Fan et al., 2024b), and SCRUB (Kurmanji et al., 2024)) on CelebA (Liu et al., 2015), Waterbirds (Sagawa et al., 2020), and FairFace (Karkkainen & Joo, 2021) in both unlearning efficacy and preserved group robustness.111We use robustness and group-robustness interchangeably in this paper.
@@ 2 Related work
Machine unlearning. Exact machine unlearning methods (Bourtoule et al., 2021; Yan et al., 2022; Aldaghri et al., 2021) guarantee that sensitive data is removed. However, they are impractical (Chundawat et al., 2023a; Kurmanji et al., 2024) as retraining (part of) the model is prohibitive (Nguyen et al., 2022). Instead, approximate unlearning concentrates on computational feasibility by relaxing the guarantees constraints (Kurmanji et al., 2024; Chundawat et al., 2023a; Jia et al., 2023; Fan et al., 2024b; Chen et al., 2023). Most works focus on random (Jia et al., 2023; Fan et al., 2024b; Cheng & Amiri, 2023; Shen et al., 2024; He et al., 2024; Zhao et al., 2024b; Huang et al., 2024), and class (Chen et al., 2023; Chundawat et al., 2023a; b; Hoang et al., 2024; Cheng et al., 2024; Zhao et al., 2024a) unlearning, respectively forgetting i.i.d. data points, and all data points from a single class. While the former lowers the forget-set accuracy to match that of the model retrained without the unlearning data, the latter aims at scoring zero accuracy on the unlearned class. A few works (Chundawat et al., 2023a; Foster et al., 2024; Cheng et al., 2024) explore subclass unlearning, where a subset of a class (e.g., subclass car from superclass vehicles) is removed from model weights. While subclass and group-robust unlearning share similarities, the latter focuses on preserving group accuracies, whereas subclass unlearning alters the model to behave as if the target subclass was never in the training set. Related to our research, Fan et al. (2024a) suggests a bi-level optimization to sample adversarial forget sets that are difficult to unlearn. Also related, Chen et al. (2024) proposes machine unlearning as an efficient debiasing technique. Instead, Zhang et al. (2024) investigates how uniform and non-uniform data sampling affect fairness in MLPs and tabular data, limiting the evaluation to exact unlearning.
Group-robust unlearning can be seen as a generalized case of random unlearning, where we drop the i.i.d. assumption of the forget set. This has implications for the design of proposed methodologies. As existing approaches assume a uniform distribution of the forget set during evaluation, they do not account for group imbalance in the forget distribution, possibly degrading the model’s robustness.
@@
Group-robust learning. Methods for group-robust optimization train deep learning models to be robust to spurious correlations. Algorithms are categorized based on their access to group information. Within those that assume access to group annotations (Sagawa et al., 2020; Goel et al., 2020; Zhang et al., 2021; Idrissi et al., 2022; Kirichenko et al., 2023), group-DRO (Sagawa et al., 2020) dynamically reweights the misclassification penalty for each group to optimize the worst-group accuracy. Instead, Idrissi et al. (2022) proposes a simple baseline that subsamples each group to match the size of the smallest one. While these works usually achieve better results, accessing the group information can be challenging (e.g., annotation cost). Within methods agnostic to group information (Liu et al., 2021; Zhang et al., 2022; Idrissi et al., 2022; Sohoni et al., 2020; Nam et al., 2020), JTT (Liu et al., 2021) increases the sampling probability of wrongly classified data to improve worst-group accuracy. Correct-N-Contrast (Zhang et al., 2022) uses a contrastive loss to pull correctly and misclassified samples of the same class while pushing apart wrongly classified data points of different categories. However, methods that require group information also show state-of-the-art performance on groups discovered from data (Kim et al., 2024; D’Incà et al., 2024).
@@ 3 Method
This section formulates the machine unlearning task (Sec.˜3.1) and the group-robust machine unlearning problem (Sec.˜3.2). Then, it introduces the proposed sample distribution reweighting strategy (Sec.˜3.3), showing its effectiveness in group-robust machine unlearning. Finally, Sec.˜3.4 describes MIU, our approximate unlearning method tailored for group-robust unlearning.
@@ 3.1 Machine unlearning
Let be a learnable function (or model), where is a non-linear feature extractor parameterized by , and is a linear classifier parameterized by , mapping inputs from the image to the target space . Let be a training dataset of size , where is an image, and its target label (e.g., age). Let a model trained on with algorithm be denoted as (or Pretrain). A machine unlearning algorithm scrubs the influence of a desired forget set from the pretrained model by outputting scrubbed weights , , such that is as close as possible to the exact unlearning model (or Retrain), trained solely on the remaining set with algorithm (Kurmanji et al., 2024; Fan et al., 2024b; Zhao et al., 2024b).
@@ 3.2 Group-robust machine unlearning
In group-robust machine unlearning, we consider the forget data non-uniformly distributed. Therefore, let us re-define the training dataset as , where and are defined as in Sec.˜3.1, and is the protected or sensitive attribute (e.g., gender, ethnicity). Now, let be the set of all groups, defined as the cartesian product between the target label set and the protected attribute set (Sagawa et al., 2020; Kirichenko et al., 2023). We denote the -th datapoint group as . Each target-sensitive attribute pair (e.g., males, between the ages of 20-29) identifies a unique group.222Target and protected attributes are color-coded.
If the forget set is non-i.i.d., the group distribution in the remaining set changes compared to the original training set. As the proportion of a group in the remaining set lowers, its resulting accuracy also decreases, which can harm the model’s generalization performance on the dominant group of the forget set. Figure˜2 shows the accuracy degradation caused by jointly removing different percentages of the least and most represented groups, i.e., blond males and non-blonde females. While the most represented group is mildly affected as its ratio drops from 44% to 7% (over four groups), the least represented one (dropping from 0.8% to 0.1%) suffers a substantial accuracy degradation of 20%. Further analysis is provided in Appx.˜B.
Therefore, the objective of group-robust machine unlearning differs from that of machine unlearning. While the latter minimizes the discrepancy with the retrained model (gold standard), the former also preserves the original model’s performance on each group.
@@ 3.3 Frustratingly easy group-robust unlearning
Section˜3.2 introduces the group-robust machine unlearning task and shows the extent to which the accuracy of the forget set dominant group(s) drops after unlearning. Unlike prior works (see Sec.˜1), we aim at unlearning non-uniform forget data while preserving the model accuracy on the dominant group of the forget set. Therefore, this section proposes an exact unlearning strategy tailored to this task. To mitigate the performance degradation of the dominant group of , we argue that reweighting the data distribution (RWG in Idrissi et al. (2022)) to account for the removed samples is a simple and effective baseline that retrains the model with a minimal performance drop. Contrary to RWG (Idrissi et al., 2022), which enforces uniform sampling distribution from group-perspective, we suggest instead to reweight the sampling distribution to match the original dataset statistics. Intuitively, increasing the sampling likelihood for partly unlearned groups rebalances the remaining set group statistics to match those of the training dataset, recovering the original robustness.
Formally, let be the probability of sampling , let respectively be the group frequencies of the training and remaining datasets. We reweight according to the ratio : , where is the adjusted group weight for sample . We denote the data distribution reweighting strategy (or reweight) as , i.e., denotes the reweighting strategy applied to the remaining dataset.
We validate reweight by comparing the model retrained with reweight and group-DRO (Sagawa et al., 2020). Figure˜3 summarizes the outcome of our analysis. As a byproduct of strongly optimizing worst-group accuracies, we notice that group-DRO (Sagawa et al., 2020) can also increase the forget set accuracy. This issue makes approximate unlearning evaluation more difficult if Retrain + group-DRO is used as the gold standard. Assuming a hypothetical original forget-set accuracy of 70%, if an approximate unlearning algorithm targeting such a gold standard leads to higher accuracy (e.g., 80%), then was the knowledge unlearned if forget-set accuracy increased? Answering this question is non-trivial, as the retrained model cannot be accessed for real applications. Furthermore, retraining with group-DRO (Sagawa et al., 2020) causes a test accuracy drop (-5.8%) in FairFace (Karkkainen & Joo, 2021). On the contrary, reweight does not suffer from these issues.
@@ 3.4 Group-robust unlearning without retraining
Section˜3.3 shows that sample reweighting is a simple and effective approach to machine unlearning that preserves the original model’s robustness. However, model retraining is inefficient (Chundawat et al., 2023a; Kurmanji et al., 2024) and unpractical (Nguyen et al., 2022). Therefore, this section proposes MIU, our approximate machine unlearning algorithm that jointly tackles unlearning while decorrelating unlearning and group robustness without retraining the entire model.
As scrubbing unlearning data affects the forget set dominant group accuracy, we propose a unified objective for jointly unlearning while preserving original group robustness. We use mutual information between output features and group annotation, which, upon minimization on forget data, jointly unlearns and mitigates the performance degradation on the dominant group of the forget set (Ragonesi et al., 2021). Formally, let be the mutual information between random variables and , associated with model features and group , then:
where is the joint pdf of and , and , are marginal pdfs of and . Minimizing reduces the dependency between and while also reducing it between and , jointly unlearning the network and disentangling features and protected attributes. However, computing the mutual information for continuous variables is generally intractable (Paninski, 2003; Belghazi et al., 2018). Therefore, we follow Belghazi et al. (2018), and estimate the mutual information with an MLP :
where is a two-layer MLP with ReLU as non-linear activation, outputting a single scalar, refers to a sampling from the joint distribution, refers to a sampling from the product of the marginal distributions. To sample from the product of the marginal distributions we sample twice from the joint distribution: , and keep only and as in Belghazi et al. (2018). We now use this definition to derive MIU (Mutual Information-aware Machine Unlearning), our proposed method.
@@
This term is high when the forget features correlate with the group information. Intuitively, we want this term to be low as this implies that we are unlearning the relation between forget-set features and labels while decorrelating the unlearning process from group information. Formally, minimizing this term achieves unlearning while preserving group robustness due to the following proposition:
Proposition 1.
Mutual information minimization is equivalent to cross-entropy loss maximization on a group classifier.
Proof.
Let us rewrite Eq.˜1 in terms of entropy and conditional entropy:
where group entropy is constant, thus, it can be ignored during optimization. Let be a group classifier, namely a linear layer that classifies image groups using features , and let be the random variable associated with estimated groups. We can now write the cross-entropy loss for classifying groups using entropy notation as:
Following Boudiaf et al. (2020), Lemma 2, we can relate Eqs.˜5 and 4 by decoupling Eq.˜5 into two steps, i.e., first optimize and then . In other words, the first step fixes the backbone weights and optimizes the group classifier weights , while the second only optimizes model weights , keeping the classifier weights frozen. As the classifier is fixed in the latter, then maximizing the cross-entropy with respect to encoder weights corresponds to maximizing , i.e., minimizing . Thus, minimizing (as in Eq.˜3) is equivalent to computing gradient ascent on group information, functionally disentangling label and spurious attributes from image features. ∎
Calibration term. As unlearning might also affect other groups, we designed an extra term to improve group performance retention. Thus, we minimize the mutual information discrepancy between the original and unlearning model on the reweighted remaining dataset :
where estimates the mutual information using features computed by the original backbone . Equation˜6 encourages the unlearned model to mimic the original robustness, preserving its behavior across groups.
Retaining term. Since unlearning generally causes performance degradation (Kurmanji et al., 2024; Fan et al., 2024b; Jia et al., 2023), we perform gradient descent steps on the remaining data to ensure that the original model group accuracy is preserved after unlearning. Additionally, we use reweight to limit the degradation of group robustness. Therefore, we optimize the cross-entropy loss on the reweighted remaining dataset :
@@
MIU pseudocode. Algorithm˜1 presents MIU pseudocode in a PyTorch-like (Paszke et al., 2019) style. We follow previous works (Kurmanji et al., 2024; Fan et al., 2024b) and compute unlearning and retaining steps separately. Thus, we alternate between computing an unlearning epoch using Eq.˜3 and a retaining one using Eqs.˜7 and 6. Like SCRUB (Kurmanji et al., 2024), we find it beneficial to stop performing unlearning steps after a predefined number of epochs (i.e., 5 out of 10). Contrary to Ragonesi et al. (2021), we do not update the mutual information representation at every step. Instead, we empirically observed that updating it via 100 gradient updates for the first epoch and 10 updates for the remaining iterations is sufficient to achieve satisfactory results and limit the required resources. As we keep the optimization steps fixed, the overall mutual information estimation overhead depends on the dataset size. For a small dataset like Waterbirds (Sagawa et al., 2020), we estimate an overhead of about . Instead, for CelebA (Liu et al., 2015) and FairFace (Karkkainen & Joo, 2021), we estimate an increase of approximately in unlearning time, which is negligible.